OpenStudy (wintersuntime):

Can someone be nice enough To help me with these problems Simplify The problems below X^2+51=100 2X^2-7X+9=0

OpenStudy (wintersuntime):

@Zale101

OpenStudy (wintersuntime):

@loser66

OpenStudy (campbell_st):

for the 1st question subtract 51 from both sides of the equation then take the square root of the answer... there are 2 solutions, the positive and negative versions.. 2nd question needs the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you have a = 2, b = -7 and c = 9 substitute them and calculate the answers

OpenStudy (wintersuntime):

okay thank you so much

OpenStudy (wintersuntime):

so for the first one would it equal 7 squared ? @campbell_st

OpenStudy (campbell_st):

well yes but the answers are x = -7 as (-7)^2 = 49 and x = 7

OpenStudy (wintersuntime):

oh okay thank you

OpenStudy (wintersuntime):

For the second one i got \[-7\pm \sqrt{-23} \over 4\]

OpenStudy (campbell_st):

that looks correct, you make need to use the complex number \[i^2 = -1\] and dix the leading term so its \[x = \frac{7 \pm \sqrt{23i^2}}{4} = \frac{7 \pm i \sqrt{23}}{4}\]

OpenStudy (wintersuntime):

why do we have to use the complex number ?

OpenStudy (campbell_st):

because you can't take the square root of a negative.... if you do it on a calculator you get math error... or some other error message.

OpenStudy (wintersuntime):

oh alright thank you, now I understand it much better.